OpenStudy (anonymous):

Solve the absolute value equation 2|3x-5|-8=8.

OpenStudy (anonymous):

Here you have two equations because of the two conditions: 1) 3x-5 >= 0 . Then equation becomes: 2*(3x-5)-8 = 8 6x-10-8= 8 6x = 26 --> x=13/3 2) 3x-5<0 ---> |3x-5| = -(3x-5) = 5-3x. Then equation becomes: 2*(5-3x)-8 = 8 10-6x-8=8 -6 = 6x ---> x = -1 x = 13/3 and x= -1 are the solutions

OpenStudy (anonymous):

Thank you!